Show that, if the chord joining the points P(ap2 , 2ap), Q(aq2 , 2aq) on the parabola y2= 4ax passes through (a, 0), then pq = −1. Further, the tangent at P meets the line through Q parallel to the axis of the parabola at R. Prove that the line x + a = 0 bisects PR.


1
Expert's answer
2021-08-10T11:35:10-0400

 The chord joining the points  "P(ap^2, 2ap)" and "Q(aq^2, 2aq)"


"y=mx+b"


"slope=m=\dfrac{2aq-2ap}{aq^2-ap^2}=\dfrac{2}{q+p}, a\not=0,q\not=\pm p,"

"y=\dfrac{2}{q+p}x+b"

Point  "P(ap^2, 2ap)"


"2ap=\dfrac{2}{q+p}(ap^2)+b"

"b=\dfrac{2(apq+ap^2-ap^2)}{q+p}"

"b=\dfrac{2apq}{q+p}"

"y=\dfrac{2}{q+p}x+\dfrac{2apq}{q+p}"

Point "(a, 0)"


"0=\dfrac{2}{q+p}a+\dfrac{2apq}{q+p}"

"pq=-1"

Therefore if the chord joining the points P(ap2 , 2ap), Q(aq2 , 2aq) on the parabola y2= 4ax passes through (a, 0), then pq = −1.

The tangent at "P(ap^2, 2ap)"


"y^2= 4ax"

"2yy'=4a"

"y'=\dfrac{2a}{y}"

"slope=\dfrac{2a}{2ap}=\dfrac{1}{p}"

"y-2ap=\dfrac{1}{p}(x-ap^2)"

"y=\dfrac{1}{p}x+ap"

The tangent at P meets the line through Q parallel to the axis of the parabola at R


"y=2aq"

"\dfrac{1}{p}x+ap=2aq"

"x=2apq-ap^2"

"R(2apq-ap^2, 2aq)"


"PR:\dfrac{x_1+x_2}{2}=\dfrac{ap^2+2apq-ap^2}{2}=apq"

"\dfrac{y_1+y_2}{2}=\dfrac{2ap+2aq}{2}=ap+aq"

The middle of "PR" is the point "(apq, ap+aq)."

If "pq = \u22121," then


"\dfrac{x_1+x_2}{2}=apq=-a"

The middle of "PR" is the point "(-a, ap+aq)."

Then the line "x+a=0" bisects PR.